Write a general formula to describe the variation
y varies directly with x
y=4 when x=20
y=?

Answer by stanbon(75887)

Write a general formula to describe the variation
y varies directly with x
y = k*x
-------------------
y=4 when x=20
k =
4 = k*20
k = 4/20
k = 1/5
-----------
Formula:
y = (x/5)
------------
cheers,
Stan H.
------